OpenStudy (wintersuntime):

How can I make a roller coaster in desmos by combining 3 equations

hero (hero):

I don't know how to create one with three equations, but here's one where you can drag the \(a_0\) term to roll the coaster along its track: https://www.desmos.com/calculator/0sgec8qzzs

OpenStudy (wintersuntime):

okay thanks and this is for my integrated 3 math class

OpenStudy (wintersuntime):

I tried that but It just hows a parabola @Mehek14

OpenStudy (l):

In general, it could be anything with a negative coefficient and only one maximum. So try something like \(f(x) = -(x-1)^4 + 100\)

OpenStudy (l):

Or to make the ride interesting and curvy, why have only one maxima? You could go with more ups and downs with\[f(x) = (x-1)(x-2)(x-3)\]

OpenStudy (wintersuntime):

okay thanks, do you know how to make a loop?

OpenStudy (l):

What do you mean by making a loop?

OpenStudy (l):

Oh, I see what you mean!

OpenStudy (wintersuntime):

yea

OpenStudy (l):

Do you want it perfectly circular? Or a little elliptical?

OpenStudy (wintersuntime):

which one do you think would be better

OpenStudy (l):

This could get weirder, but the equation of a circle is\[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (wintersuntime):

ok

OpenStudy (supercalifragalisticexspeaalli):

lol u still need some help